OpenStudy (anonymous):

whats the slope of 23x-10y=-800

OpenStudy (anonymous):

the slope of 23x-10y=800 is you have to simplify the equation add 23x to both sides of the equal sign -10y=23x-800 then you divid both sides by -10 y=23/10x-800/10 y= 23/10x-80

OpenStudy (anonymous):

slope is always mx which is y=mx+b